# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\eta},g}$
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\alpha(g_{0})^{-1}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (98)
@@
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0})|s_{d,0})=\alpha^{-1}(g_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\left[(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})-2m^{-1}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0})^{T}\right]\mathbf{D}_{m}^{\#^{T}}\underline{\mathbf{I}}_{m}^{T}.$ (99)
@@ V-B Equality chains for parametric and semiparametric information bounds in RES distributions
We are finally ready to introduce two sequences of equalities among information bounds for the parameters of interest in RES distributions. The following proposition summarizes the key points from the previous sections and can serve as a take-away message for readers who are not interested in full mathematical details.
@@
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ be a RES distributed vector with location vector $\bm{\mu}_{0}$, scatter matrix $\bm{\Sigma}_{0}$ and density generator $g_{0}\in\bar{\mathcal{G}}$. Let $S(\bm{\Sigma}_{0})=s_{0}$ be a constraint on the covariance matrix where $S$ is a matrix function satisfying Assumptions A1, A2 and A3 and let $\mathbf{V}_{S,0}=s_{0}^{-1}\bm{\Sigma}_{0}$ the related shape matrix. Then, the following chain of equalities hold:
@@
    Relation among the SCRB for the location vector $\bm{\mu}_{0}$ in the semiparametric model $\mathcal{P}_{\bm{\eta},g}$ and three different CRB in the parametric models $\mathcal{P}_{\bm{\eta},s}$, $\mathcal{P}_{\bm{\eta},s,\bm{\zeta}}$ and $\mathcal{P}_{\bm{\eta}}$:
@@
    Relation among the SCRB for $\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})$ in the canonical semiparametric model $\mathcal{P}_{\bm{\eta},g}$ and three different parametric CRB (CRB) in the parametric models $\mathcal{P}_{\bm{\eta},s}$, $\mathcal{P}_{\bm{\eta},s,\bm{\zeta}}$ and $\mathcal{P}_{\bm{\eta}}$:
[…]
    where $\mathbf{P}_{S}(\mathbf{V}_{S,0})\triangleq\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0})\bm{\nabla}_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})$.
Proof:
The proof of the Proposition V.2 follow directly from the Propositions IV.1 and IV.2 and in particular from the fact that:
@@
    $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$,
@@
    If $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then $\bar{\mathbf{I}}(\bm{\gamma}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$.
@@ VI-A Some preliminaries
Let $\mathcal{X}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g)$ be a RES-distributed random vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})\in\mathbb{R}^{m}$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})\in\mathcal{S}_{m}^{\mathbb{R}}$ are parameterized by a $d$-dimensional parameter vector $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Theta\triangleq\Gamma\times\Psi$. As in Sect. II, $\Gamma\subseteq\mathbb{R}^{q}$ denotes the set of the (finite-dimensional) parameter vectors $\bm{\gamma}$ of interests, $\Psi\subseteq\mathbb{R}^{r}$ denotes the set of (finite-dimensional) nuisance parameter vectors $\bm{\xi}$ and $\bar{\mathcal{G}}$ is the set of the (infinite-dimensional) nuisance functions $g$ already defined in (34). In the following, we always assume that the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ satisfies the following assumptions:
@@
    the inverse $\bm{\Sigma}(\bm{\theta})]^{-1}$ exists for all $\bm{\theta}\in\Theta$.
The related semiparametric model can be then expressed as:
$\mathcal{P}_{\bm{\theta},g}=\left\{p_{X}(\mathbf{x}|,\bm{\theta},g)=|\bm{\Sigma}(\bm{\theta})|^{-1/2}g\left((\mathbf{x}-\bm{\mu}(\bm{\theta}))^{T}\bm{\Sigma}(\bm{\theta})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\theta}))\right);\bm{\theta}\in\Theta,g\in\overline{\mathcal{G}}\right\},$ (102)
where, from (32), we have that:
$(\mathbf{x}-\bm{\mu}(\bm{\theta}_{0}))^{T}\bm{\Sigma}(\bm{\theta}_{0})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\theta}_{0}))\triangleq Q_{0}=_{d}\mathcal{Q}.$ (103)
In line with our earlier analysis, we may pose the following question: is it possible to characterize all parameterisations $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ that imply that not knowing the finite-dimensional nuisance vector $\bm{\xi}_{0}\in\Psi$, while knowing $g_{0}$, leads to the same loss of efficiency as not knowing $\bm{\xi}_{0}\in\Psi$ and the density generator $g_{0}\in\overline{\mathcal{G}}$ (i.e. the infinite-dimensional nuisance parameter)?
To answer this question, let us start by evaluating the score vector
$\mathbf{s}_{\bm{\theta}_{0}}=(\mathbf{s}_{\bm{\gamma}_{0}}^{T},\;\mathbf{s}_{\bm{\xi}_{0}}^{T})^{T},$ (104)
and the related FIM $\mathbf{I}_{\bm{\theta}_{0}}={E_{0}\{\mathbf{s}_{\bm{\theta}_{0}}\mathbf{s}_{\bm{\theta}_{0}}^{T}\}}$ of the parametric model $\mathcal{P}_{\bm{\theta},g_{0}}$. From the definition of the parametric and non-parametric submodels of a semiparametric model given in Sect. II, it is immediate to verify that $\mathcal{P}_{\bm{\theta},g_{0}}$ correspond to the parametric submodel $\mathcal{P}_{1,2}$ in (8). Moreover, as a direct consequence of (103), we have that the non-parametric submodel $\mathcal{P}_{g}$ of $\mathcal{P}_{\bm{\theta},g}$ (that correspond to $\mathcal{P}_{3}$ in (9)) of $\mathcal{P}_{\bm{\theta},g}$ admits the same tangent space $\mathcal{T}_{g_{0}}$ already derived in (75). As the consequence, the orthogonal projection of a generic element $h\in\mathcal{H}$ onto $\mathcal{T}_{g_{0}}$ is again the conditional expectation given in (76).
Let us start with the calculation of the score function $\mathbf{s}_{\bm{\theta}_{0}}$. Following the derivation in [26, Sec. 3.1] and [27, Sec. III], each entry of $\mathbf{s}_{\bm{\theta}_{0}}$ can be easily evaluated as:
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}&\triangleq\left.\frac{\partial\ln p_{X}\left(\mathbf{x};\bm{\theta},g_{0}\right)}{\partial\theta_{i}}\right|_{\bm{\theta}=\bm{\theta}_{0}}=-2^{-1}\left(\mathrm{tr}(\mathbf{P}_{i}^{0})+\varphi_{0}(Q_{0})Q_{i}^{0}\right),\end{split}$ (105)
where, according to the adopted notation $Q_{i}^{0}\triangleq\frac{\partial Q_{0}}{\partial\theta_{i}}$, the function $\varphi_{0}$ has already been defined in (58), and
$\mathbf{P}_{i}^{0}\triangleq\bm{\Sigma}_{0}^{-1/2}\bm{\Sigma}_{i}^{0}\bm{\Sigma}_{0}^{-1/2}.$ (106)
Moreover, from direct calculation, we have:
$Q_{i}^{0}=-2(\mathbf{x}-\bm{\mu}_{0})^{T}\bm{\Sigma}_{0}^{-1}\bm{\mu}_{i}^{0}-(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{S}_{i}^{0}(\mathbf{x}-\bm{\mu}_{0}),$ (107)
where $\bm{\mu}_{i}^{0}\triangleq\frac{\partial\bm{\mu}_{0}}{\partial\theta_{i}}$ and
$\mathbf{S}_{i}^{0}=\bm{\Sigma}_{0}^{-1}\bm{\Sigma}_{i}^{0}\bm{\Sigma}_{0}^{-1}.$ (108)
By collecting previous results, the entries of the score vector $\mathbf{s}_{\bm{\theta}_{0}}$ can be expressed as:
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}=-2^{-1}\mathrm{tr}(\mathbf{P}_{i}^{0})+\varphi_{0}(Q_{0})(\mathbf{x}-\bm{\mu}_{0})^{T}\bm{\Sigma}_{0}^{-1}\bm{\mu}_{i}^{0}+2^{-1}\varphi_{0}(Q_{0})(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{S}_{i}^{0}(\mathbf{x}-\bm{\mu}_{0}),\;i=1,\ldots,d.\end{split}$ (109)
Using the stochastic representation in (31), the score vector in (109) can be rewritten as:

### v3

@@ III-A Essentials on RES distributions
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\alpha(\bar{g}_{0})^{-1}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (99)
@@
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0})|s_{d,0})=\alpha(\bar{g}_{0})^{-1}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\left[(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})-2m^{-1}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0})^{T}\right]\mathbf{D}_{m}^{\#^{T}}\underline{\mathbf{I}}_{m}^{T}.$ (100)
@@ V-B Equality chains for parametric and semiparametric information bounds in RES distributions
The following proposition summarizes the key points from the previous sections and can serve as a take-away message for readers who are not interested in full mathematical details.
@@
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},\bar{g}_{0})=RES_{m}(\bm{\mu}_{0},\mathbf{V}_{S,0},g_{0})$ be a RES distributed vector with location vector $\bm{\mu}_{0}$, scatter matrix $\bm{\Sigma}_{0}=s_{0}\mathbf{V}_{S,0}$ and (constrained or unconstrained) density generator $\bar{g}_{0}\in\overline{\mathcal{G}}$ or $g_{0}\in\mathcal{G}$, such that $g_{0}(t)=s_{0}^{-m/2}\bar{g}_{0}(s_{0}^{-1}t)$, $t\in\mathbb{R}^{+}$. Let $S(\bm{\Sigma}_{0})=s_{0}$ be a scale function satisfying A1, A2 and A3 and let $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$ the relevant shape matrix. Then, the following chain of equalities hold:
[…]
where $\mathbf{P}_{S}(\mathbf{V}_{S,0})\triangleq\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0})\bm{\nabla}_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})$. A similar result holds true for the location vector $\bm{\mu}_{0}$.
Proof:
The proof of the Proposition V.2 follows directly from the Propositions IV.1 and Corollary IV.1 and in particular from the fact that:
@@
    $\bar{\mathbf{I}}(\bm{\gamma}_{0}|g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$,
  2. 2.
    If $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then $\bar{\mathbf{I}}_{\bm{\gamma}_{0}}=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$.
V-C Attainability of the semiparametric bound: the $R$-estimator
A natural question at this point concerns the attainability of the previously derived semiparametric bound on the shape matrix $\mathbf{V}_{S,0}$. Indeed, to address this question, in the seminal work [28], a rank-based $R$-estimator for the shape matrix has been proposed and its asymptotic performance derived. In their work, the authors consider the scale function $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$. In this subsection, we will show how to construct this estimator for any scale function $S$. This generalization follows directly from the outcomes of the recent work [29] which provides a general framework on constrained information bounds in semiparametric models. Due to space limitations, we will not discuss here the properties or theoretical basis of this estimator. Interested readers are referred to [12, 28, 10, 29] while a tutorial introduction more in line with the approach adopted in this paper can be found in [30, 31]. According to the findings presented in the previous section, without loss of generality we assume that $\bm{\mu}={\bf 0}$.
Let $\mathbf{x}_{l}\sim RES_{m}(\mathbf{0},\mathbf{V}_{S,0},g_{0})$, $l=1,\ldots,n$ be a set of independent, identically distributed (iid), RES distributed, observations. The semiparametric $R$-estimator of $\mathbf{V}_{S,0}$ is given by:
$\mathrm{vecs}(\widehat{\mathbf{V}}_{S,R})=\mathrm{vecs}(\widehat{\mathbf{V}}_{S}^{\star})+\frac{1}{\hat{\alpha}\sqrt{n}}\bm{\Xi}_{\widehat{\mathbf{V}}_{S}^{\star}}\bm{\Delta}_{\widehat{\mathbf{V}}_{S}^{\star}},$ (102)
where $\widehat{\mathbf{V}}_{S}^{\star}$ is a preliminary, $\sqrt{n}$-consistent, estimator of $\mathbf{V}_{S,0}$. The matrix $\bm{\Xi}_{\widehat{\mathbf{V}}_{S}^{\star}}$ and the vector $\bm{\Delta}_{\widehat{\mathbf{V}}_{S}^{\star}}$ are defined as:
$\bm{\Xi}_{\widehat{\mathbf{V}}_{S}^{\star}}\triangleq 2\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}\left[\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}^{T}\bm{\Upsilon}_{\widehat{\mathbf{V}}_{S}^{\star}}\bm{\Upsilon}_{\widehat{\mathbf{V}}_{S}^{\star}}^{T}\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}\right]^{-1}\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}^{T},$ (103)
$\bm{\Delta}_{\widehat{\mathbf{V}}_{S}^{\star}}\triangleq\frac{1}{2\sqrt{n}}\bm{\Upsilon}_{\widehat{\mathbf{V}}_{S}^{\star}}\sum_{l=1}^{n}K\left(\frac{r^{\star}_{l}}{n+1}\right)\mathrm{vec}(\hat{\mathbf{u}}^{\star}_{l}(\hat{\mathbf{u}}^{\star}_{l})^{T}),$ (104)
and the scalar $\hat{\alpha}$ can be obtained as [28, Sect. 4].
The function $K:(0,1)\rightarrow\mathbb{R}^{+}$ is generally called the score function and belongs to the set $\mathcal{K}$ of continuous, square integrable functions that can be expressed as the difference of two monotone increasing functions [28, 32].
To define the $R$-estimator in (102), we also need the following terms:
  • $\hat{Q}^{\star}_{l}\triangleq\mathbf{x}_{l}^{T}[\widehat{\mathbf{V}}^{\star}_{S}]^{-1}\mathbf{x}_{l}$,
  • $\hat{\mathbf{u}}^{\star}_{l}\triangleq(\hat{Q}^{\star}_{l})^{-1/2}[\widehat{\mathbf{V}}^{\star}_{S}]^{-1/2}\mathbf{x}_{l}$,
  • $\{r^{\star}_{l}\}_{l=1}^{n}$ are the ranks of the (continuous) real random variables $\{\hat{Q}^{\star}_{l}\}_{l=1}^{n}$,
  • $\bm{\Upsilon}_{\widehat{\mathbf{V}}^{\star}_{S}}\triangleq\mathbf{D}_{m}^{T}\left([\widehat{\mathbf{V}}^{\star}_{S}]^{-1/2}\otimes[\widehat{\mathbf{V}}^{\star}_{S}]^{-1/2}\right)\Pi^{\perp}_{\mathrm{vec}(\mathbf{I}_{m})}$, where $\Pi^{\perp}_{\mathrm{vec}(\mathbf{I}_{m})}=\mathbf{I}_{m^{2}}-m^{-1}\mathrm{vec}(\mathbf{I}_{m})\mathrm{vec}(\mathbf{I}_{m})^{T}$,
  • $\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}$ is the matrix defined in Appendix .7 in (177)-(179).
To show the effectiveness, in terms of efficiency and robustness, of the $R$-estimator in (102), we conducted a simulation study for finite $n$. Specifically, we generate $n$ iid observation vectors drawn form a $t$-distribution with $\nu\in(2,+\infty)$ degrees of freedom [9, Sect. 1.5.2] and we evaluated the following Mean Squared Error (MSE) index:
@@ V-C Attainability of the semiparametric bound: the $R$-estimator
for the following three estimators:
  1. 1.
    the constrained Sample Covariance Matrix $\widehat{\mathbf{V}}_{S,SCM}$,
@@
    the constrained Tyler estimator $\widehat{\mathbf{V}}_{S,Tyler}$ [33],
  3. 3.
    the $R$-estimator in (102) where the Tyler estimator is used as preliminary estimator, i.e. $\widehat{\mathbf{V}}_{S}^{\star}=\widehat{\mathbf{V}}_{S,Tyler}$. As score function $K$ we used:
    • the van der Waerden score $K_{vdW}(u)\triangleq\Psi^{-1}(u)$ where $\Psi^{-1}$ indicates the inverse function of the cdf of a central chi-squared-distributed random variable with $m$ degrees of freedom [12]. As discussed in [28, 30], $K_{vdW}$ has been proved to provide a good compromise between robustness and efficiency. Moreover, the resulting $R$-estimator is asymptotically equivalent to the (constrained) maximum likelihood estimator of the shape matrix for Gaussian data.
    • the $t_{\nu}$-based score $K_{t_{\nu}}(u)=\frac{m(m+\nu)F^{-1}_{m,\nu}(u)}{(\nu+mF^{-1}_{m,\nu}(u))}$ where $F_{m,\nu}(u)$ stands for the cdf of a Fisher random variable with $m$ and $\nu\in(0,\infty)$ degrees of freedom [12]. Here the parameter $\nu$ should be thought of as a knob to adjust the trade-off between robustness and efficiency (at Gaussian data). In fact, small $\nu$ values favor robustness while large values promote efficiency (at Gaussian data), and in particular $\lim_{\nu\rightarrow\infty}K_{t_{\nu}}(u)=K_{vdW}(u)$. Also in this case, the resulting $R$-estimator is asymptotically equivalent to the (constrained) maximum likelihood estimator of the shape matrix for $t_{\nu}$-distributed data.
The MSE of the above mentioned estimators is compared to the trace of the semiparametric CRB for $\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})$ obtained in (101) and of the parametric CRB (numerically derived from the inversion of $\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ in (69)) for the estimation of $\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})$ in the parametric sub-model $\mathcal{P}_{\bm{\gamma}}$ in (43), i.e. when the scale $s$ and the density generator are perfectly known.
In our simulations we generated $n=100$ $t$-distributed observation vectors of dimension $m=4$ with a degrees of freedom $\nu$ varying from $2.1$ to $20$. For the $t$-distribution, the coefficient in (63) is equal to $\alpha(\bar{g}_{0})=\frac{m+\nu}{2+m+\nu}$. The scatter matrix used to generate the data has the Toeplitz form $\bm{\Sigma}_{0}]_{ij}=\rho_{0}^{|i-j|}$ with $\rho_{0}=0.8$. The MSE indices of the estimators have been evaluated by means of $10^{5}$ Monte-Carlo trials. A preliminary version of the code is available in [34]. The simulation analysis is conducted for the three scale functions: $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$ (Fig. 1), $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ (Fig. 2) and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ (Fig. 3).
The simulations show that the $R$-estimator $\widehat{\mathbf{V}}_{S,R}$ presents better performance in terms of efficiency and robustness, compared to $\widehat{\mathbf{V}}_{S,SCM}$ (neither robust nor efficient) and $\widehat{\mathbf{V}}_{S,Tyler}$ (robust but not efficient). We note, however, that as $\nu\rightarrow\infty$,i.e., as the distribution tends to Gaussian, the $\widehat{\mathbf{V}}_{S,SCM}$ attains the lowest MSE, it is then the optimal estimator. As expected, one can note that the $K_{t_{4}}$ score function provides better performance, with respect to the van der Waerden score $K_{vdW}$ for heavy-tailed data ($\nu\approx 2$). On the other hand, $K_{vdW}$ is preferable when the data are closed to Gaussian ($\nu\gg 2$). Moreover, Fig. 3 confirms that, as expected from Proposition V.2, when the determinant-based scale is adopted, the semiparametric CRB is equal to the parametric $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0}))$ in $\mathcal{P}_{\bm{\gamma}}$ in (43).
A comment on the implementation of the $R$-estimator in (102) is in order. Since, in general, the matrix $\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}$ depends on the preliminary estimator, it may be of interest to iterate the estimation process. Specifically, as preliminary estimator at step $k$, we may use the $\sqrt{n}$-consistent $R$-estimate obtained at the step k1k-1. For example, we may implement the following iterations:
$\left\{\begin{array}[]{l}\widehat{\mathbf{V}}_{S,R}^{(0)}=\widehat{\mathbf{V}}_{S,Tyler},\\ \mathrm{vecs}(\widehat{\mathbf{V}}_{S,R}^{(k+1)})=\mathrm{vecs}(\widehat{\mathbf{V}}_{S,R}^{(k)})+\frac{1}{\hat{\alpha}\sqrt{n}}\bm{\Xi}_{\widehat{\mathbf{V}}_{S,R}^{(k)}}\bm{\Delta}_{\widehat{\mathbf{V}}_{S,R}^{(k)}},\end{array}\right$. (106)
for $k=0,\ldots,K$. While this iterative procedure does not have any impact on the efficiency of the resulting $R$-estimator, it may bring some improvement in the finite-sample regime. On the contrary, this procedure leads to no or negligible improvement when the matrix $\mathbf{U}_{\widehat{\mathbf{V}}_{S}^{\star}}$ does not depend on the preliminary estimator, as in the case of the scale functions $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$ and $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$.
Clearly, this subsection does not resolve all questions about the possibility to derive semiparametric efficient estimates of the shape matrix $\mathbf{V}_{S,0}$. Further in-depth analysis is needed, both from a theoretical and a simulation perspectives. Special attention should be devoted to the following two points:
@@
    To derive $\hat{\alpha}$ in (102), we used the $\sqrt{n}$-consistent but not efficient estimator proposed in [28, eq. (4.1)]. An efficient estimation method has been proposed in [28, Sect. 4.2], but presents a considerable computational (numerical) difficulty. Future works will then address its efficient numerical implementation.
@@
    A limitation of the estimator in (102) is that it satisfies the constraint $S(\widehat{\mathbf{V}}_{S,R})=1$ only asymptotically as $n\rightarrow\infty$ [29, Theo. 2]. A possible way to “force” the constraint for each $n$ is to derive $\widehat{\mathbf{V}}_{S,R}$ directly on the differential manifold induced by $S$, i.e. $\mathcal{S}_{m,S}^{\mathbb{R}}$ in (38). This important issue may be addressed building upon the recent and seminal work [35]. Again, this original line of research is left for future works.
@@ VI-A Some preliminaries
Let $\mathcal{X}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},\bar{g}_{0})$ be a RES-distributed random vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})\in\mathbb{R}^{m}$ and scatter matrix (which, under the constraint (34), equates the covariance of $\mathbf{x}$ as shown in (35)) $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})\in\mathcal{S}_{m}^{\mathbb{R}}$ are parameterized by a $d$-dimensional parameter vector $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Theta\triangleq\Gamma\times\Psi$. As in Sect. II, $\Gamma\subseteq\mathbb{R}^{q}$ denotes the set of the (finite-dimensional) parameter vectors $\bm{\gamma}$ of interests, $\Psi\subseteq\mathbb{R}^{r}$ denotes the set of (finite-dimensional) nuisance parameter vectors $\bm{\xi}$ and $g_{0}\in\mathcal{G}$ is the (infinite-dimensional) nuisance functions. In the following, we always assume that the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ satisfies the following assumptions:
@@
    the inverse $\bm{\Sigma}(\bm{\theta})^{-1}$ exists for all $\bm{\theta}\in\Theta$.
We note that, from (32), we have:
$(\mathbf{x}-\bm{\mu}(\bm{\theta}))^{T}\bm{\Sigma}(\bm{\theta})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\theta}))=Q_{\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta})}({\bf x})=_{d}\mathcal{Q}\;\forall\bm{\theta}\in\Theta.$ (107)
As in Sect. III-A, we can then build the semiparametric model:
$\mathcal{P}_{\bm{\theta},\bar{g}}=\left\{p_{X}(\mathbf{x}|\bm{\theta},\bar{g})=|\bm{\Sigma}(\bm{\theta})|^{-1/2}\bar{g}\left(Q_{\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta})}({\bf x})\right);\bm{\theta}\in\Theta,\bar{g}\in\overline{\mathcal{G}}\right\}.$ (108)
We note, in passing that, if we choose as parameter vector $\bm{\theta}=(\bm{\gamma}^{T},s)^{T}$ in (42), we immediately have that $\mathcal{P}_{\bm{\theta},\bar{g}}$ correspond to $\mathcal{P}_{\bm{\eta},\bar{g}}$ in (40).
Moreover, exactly as for the semiparametric models presented Sect. III-A, for the model $\mathcal{P}_{\bm{\theta},\bar{g}}$, we may identify the parametric sub-model:
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},\bar{g}_{0}):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\},$ (109)
and the non-parametric sub-model (that is the same of the one introduced in (46)) as:
$\mathcal{P}^{c}_{\bar{g}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},\bm{\xi}_{0},\bar{g}):\bar{g}\in\overline{\mathcal{G}}\right\}.$ (110)
Note that the (finite-dimensional) nuisance tangent space for the parametric sub-model $\mathcal{P}_{\bm{\gamma},\bm{\xi}}$ can be expressed as
$\mathcal{T}_{\bm{\xi}_{0}}=\mathrm{Span}\{[\mathbf{s}_{\bm{\xi}_{0}}]_{1},\ldots,[\mathbf{s}_{\bm{\xi}_{0}}]_{r}\},$ (111)
where $\mathbf{s}_{\bm{\xi}_{0}}$ is the score vector of the nuisance parameters $\bm{\xi}_{0}\in\Psi$. Moreover, the (infinite-dimensional) nuisance tangent space $\mathcal{T}^{c}_{\bar{g}_{0}}$ for the non-parametric model $\mathcal{P}^{c}_{\bar{g}}$, along with its relative projection operator $\Pi(h|\mathcal{T}^{c}_{\bar{g}_{0}})$, are the ones given in (79) and (81), respectively.
To compute this projection operator, let us start by evaluating the score vectors in the parametric sub-model $\mathcal{P}_{\bm{\gamma},\bm{\xi}}$ in (109), i.e. $\mathbf{s}_{\bm{\theta}_{0}}=(\mathbf{s}_{\bm{\gamma}_{0}}^{T},\;\mathbf{s}_{\bm{\xi}_{0}}^{T})^{T}$ and the related FIM $\mathbf{I}_{\bm{\theta}_{0}}={E_{0}\{\mathbf{s}_{\bm{\theta}_{0}}\mathbf{s}_{\bm{\theta}_{0}}^{T}\}}$. Following the derivation in [36, Sec. 3.1] and [37, Sec. III], each entry of $\mathbf{s}_{\bm{\theta}_{0}}$ can be expressed as: